According to a study done last year, the average monthly expenses for mobile phone loads of college students in San Mateo, Rizal was P 400.00. A statistics student believes that this amount has decreased since January of this year. Is there a reason to believe that this amount has decreased if a random sample of 50 students has an average monthly expense for mobile phone loads of P 380.00? Use .05 level of significance. Assume the population standard deviation is P 75.00.  


1
Expert's answer
2021-04-14T14:11:41-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq400"

"H_1:\mu<400"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a left-tailed test is "z_c=-1.6449."

The rejection region for this left-tailed test is "R:\{z:z<-1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{x-\mu}{\sigma/\sqrt{n}}=\dfrac{380-400}{75.00/\sqrt{50}}=-1.8856"

Since it is observed that "z=-1.8856<-1.6449=z_c,"  it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is less than 400, at the 0.05 significance level.


Using the P-value approach: The p-value is "p=P(Z<-1.8856)=0.0297," and since "p=0.0297<0.05=\alpha," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is less than 400, at the 0.05 significance level.